OpenStudy (anonymous):

Find the derivative of the function. h(x) = log3(x^2 + x) h'(x)=

OpenStudy (saifoo.khan):

is 3 the base?

OpenStudy (anonymous):

log(3)

OpenStudy (anonymous):

\[h'(x)=\frac{2 x+1}{\left(x^2+x\right) \log (3)} \]